OpenStudy (anonymous):

solve the equation 4(x-2)=10x-40 A: x= -32/6 B: x= 12/7 C: x= 6 1/3 D: x= 5 1/3

OpenStudy (anonymous):

4x - 8 = 10x -40 6x=40-8

OpenStudy (anonymous):

so do you think it would be A?

OpenStudy (anonymous):

actually it's 32/6 i guess......not -32/6

OpenStudy (anonymous):

huh ... ok THANKS !!!!

jimthompson5910 (jim_thompson5910):

32/6 = (30+2)/6 = 30/6 + 2/6 = 5 + 2/6 = 5+1/3 So \[\Large \frac{32}{6}=5\frac{1}{3}\]

OpenStudy (anonymous):

THANK YOU jim haha

jimthompson5910 (jim_thompson5910):

you're welcome